# On schurity of the dihedral group $\boldsymbol{D_{2p}}$

### v3

@@ 5. Further discussion
We do not know whether it is possible to extend Theorem 1.1 to other primes $p$. Recall that if $p\equiv 3\mod~{}4$ and $p>11$, then $D_{2p}$ is not Schur by [26, Corollary 5.3]. In this case, there is a Paley difference $D$ set in $C_{p}$ and the $S$-ring $\mathcal{A}(D)$ is nonschurian whenever $p>11$. However, $D_{2p}$ can be not Schur even if $p\equiv 1\mod~{}4$ which follows from the lemma below.
@@
From [3, Theorem VI.8.11(b)] it follows that there is a biquadric residue difference set $D$ in a cyclic group of order $p$ with parameters $(v,k,\lambda)=(4t^{2}+1,t^{2},\frac{t^{2}-1}{4})$ if $p=4t^{2}+1$ and $(v,k,\lambda)=(4t^{2}+9,t^{2}+3,\frac{t^{2}+3}{4})$ if $p=4t^{2}+9$. Suppose that $p=4t^{2}+1$. Then $k=t^{2}$ divides $v-1=4t^{2}$. If $t>3$, then $41+\sqrt{k}=1+t>\frac{v-1}{k}=$ 4. So the design corresponding to $D$ is not $2$-transitive by [18, Theorem 8.3]. Therefore the $S$-ring over $D_{2p}$ constructed from $D$ as in [26, Lemma 5.1] is nonschurian by [26, Lemma 5.2] and hence $D_{2p}$ is not Schur. If $t=3$, then it is easy to see that $v=4t^{2}+1=37$ can not be presented in the form $\frac{q^{d+1}-1}{q-1}$ for some prime power $q$ and $d\geq 2$ and hence $D_{2p}$ is not Schur by [26, Corollary 5.4].
Now let $p=4t^{2}+9$. Assume that $D_{2p}$ is Schur. Then [26, Corollary 5.4] implies that $D$ must have parameters
@@
Observe that if $p=4t^{2}+9$, then $p-1$ can be of the form $p-1=4m$, where $m$ is odd square-free. For example, $p=3373=4\cdot 29^{2}+9=4\cdot 3\cdot 281+1$. It would be interesting to study schurity of $D_{2p}$ for all primes $p$ such that $p\equiv 1\mod~{}4$.

### v4

@@ 8. Further discussion
8.1. $S$-rings from relative difference sets
As it was mentioned in the introduction, the only non-Schur dihedral groups of order at most $62$ which satisfy all the conditions from Theorem 1.2 and Corollary 6.8 are $D_{52}$ and $D_{60}$. Nonschurian $S$-rings over these groups arise from relative difference sets. Let us discuss their constructions.
Let $A$ be a group and $A_{0}$ a proper nontrivial subgroup of $A$ of order and index $n$ and $m$, respectively. A subset $D\subseteq A$ of size $k$ is called a relative difference set (RDS for short) with forbidden subgroup $A_{0}$ if
(23) $\underline{D}\cdot\underline{D}^{-1}=ke+\lambda(\underline{A}-\underline{A_{0}})$
for some positive integer $\lambda$. The numbers $(m,n,k,\lambda)$ are called the parameters of $D$. Due to the definition of RDS, we have $|D\cap A_{0}a|\leq 1$ for every $a\in A$. Suppose that $A$ is abelian, $G$ is a generalized dihedral group associated with $A$, and $b\in G\setminus A$. Consider the partition $\mathcal{S}$ of $G$ into the following sets:
$X_{a}=\{a\},~{}X_{0}=A\setminus A_{0},~{}Y_{a}=Dab,~{}Y_{0}=(A\setminus DA_{0})b,~{}a\in A_{0}.$
Lemma 8.1.
The $\mathbb{Z}$-module $\mathcal{A}=\mathcal{A}(D)=\operatorname{Span}_{\mathbb{Z}}\{\underline{X}:~{}X\in\mathcal{S}\}$ is an $S$-ring over $G$.
Proof.
One can see that every set from $\mathcal{S}$ outside $A_{0}$ is inverse-closed and hence $X^{-1}\in\mathcal{S}$ for every $X\in\mathcal{S}$. Since $Y_{0}$ is a complement to the union of all other sets from $\mathcal{S}$, it suffices to verify that $\underline{X}\cdot\underline{Y}\in\mathcal{A}$ for all $X,Y\in\{X_{a},X_{0},Y_{a}:~{}a\in A_{0}\}$. A straightforward computation using Eq. (23) in the group ring $\mathbb{Z}G$ implies that
$\underline{X_{a_{1}}}\cdot\underline{X_{a_{2}}}=\underline{X_{a_{1}a_{2}}},~{}\underline{X_{a}}\cdot\underline{X_{0}}=\underline{X_{0}}\cdot\underline{X_{a}}=\underline{X_{0}},~{}\underline{X_{a_{1}}}\cdot\underline{Y_{a_{2}}}=\underline{Y_{a_{1}a_{2}}},~{}\underline{Y_{a_{2}}}\cdot\underline{X_{a_{1}}}=\underline{Y_{a_{1}^{-1}a_{2}}},$
$\underline{X_{0}}^{2}=(mn-2n)\underline{X_{0}}+(mn-n)\underline{A_{0}},~{}\underline{X_{0}}\cdot\underline{Y_{a}}=\underline{Y_{a}}\cdot\underline{X_{0}}=k\underline{Ab}-\underline{Y_{0}},$
$\underline{Y_{a_{1}}}\cdot\underline{Y_{a_{2}}}=k\underline{X_{a_{1}a_{2}^{-1}}}+\lambda\underline{X_{0}},$
for all $a,a_{1},a_{2}\in A_{0}$. ∎
From [1, Theorem 1.2] it follows that there exists an RDS $D$ with parameters
$\left(\frac{q^{d}-1}{q-1},n,q^{d-1},\frac{q^{d-2}(q-1)}{n}\right)$
in a cyclic group, where $q$ is a prime power and $d,n\geq 2$ are integers, if and only if $n$ divides $q-1$ when $q$ is odd or $d$ is even and $n$ divides $2(q-1)$ when $q$ is even and $d$ is odd. Therefore one can construct an $S$-ring $\mathcal{A}(D)$ over a dihedral group of order $2l$ for each $l$ of the form $l=\frac{(q^{d+1}-1)n}{q-1}$, where $q$, $d$, and $n$ satisfy the above condition, in particular for $l=15$ and $l=26$. The computational results [44] imply the following: $(1)$ $\mathcal{A}(D)$ is schurian if $l<26$; $(2)$ if $l=26$, then $\mathcal{A}(D)$ is a unique nonschurian $S$-ring over $D_{52}$; $(3)$ a nonschurian $S$-ring over $D_{60}$ can be obtained as a generalized wreath product of a schurian $S$-ring $\mathcal{A}(D)$ over $D_{30}$ and some other $S$-ring. It would be interesting to study a schurity of $\mathcal{A}(D)$ in a general case. Finally, it should be mentioned that the Cayley graph $\operatorname{Cay}(G,Db)$, whose connection set is a basic set of $\mathcal{A}(D)$, is exactly a graph from [34, Proposition 5.1] appeared in the context of studying $2$-arc transitive dihedrants (see also [21]).
8.2. More on schurity of $D_{2p}$
We do not know whether it is possible to extend Theorem 1.3 to other primes $p$. Recall that if $p\equiv 3\bmod 4$ and $p>11$, then $D_{2p}$ is not Schur by Lemma 6.3. In this case, there is a Paley difference $D$ set in $C_{p}$ and the $S$-ring $\mathcal{A}(D)$ is nonschurian whenever $p>11$. However, $D_{2p}$ can be not Schur even if $p\equiv 1\bmod 4$ which follows from the lemma below.
@@ 8.2. More on schurity of $D_{2p}$
From [4, Theorem VI.8.11(b)] it follows that there is a biquadric residue difference set $D$ in a cyclic group of order $p$ with parameters $(v,k,\lambda)=(4t^{2}+1,t^{2},\frac{t^{2}-1}{4})$ if $p=4t^{2}+1$ and $(v,k,\lambda)=(4t^{2}+9,t^{2}+3,\frac{t^{2}+3}{4})$ if $p=4t^{2}+9$. Suppose that $p=4t^{2}+1$. Then $k=t^{2}$ divides $v-1=4t^{2}$. If $t>3$, then $41+\sqrt{k}=1+t>\frac{v-1}{k}=$ 4. So the design corresponding to $D$ is not $2$-transitive by [23, Theorem 8.3]. Therefore the $S$-ring $\mathcal{A}(D)$ over $D_{2p}$ constructed from $D$ as in Section $3$ is nonschurian by [32, Lemma 5.2] and hence $D_{2p}$ is not Schur. If $t=3$, then it is easy to see that $v=4t^{2}+1=37$ can not be presented in the form $\frac{q^{d+1}-1}{q-1}$ for some prime power $q$ and $d\geq 2$ and hence $D_{2p}$ is not Schur by Lemma 6.4.
Now let $p=4t^{2}+9$. Assume that $D_{2p}$ is Schur. Then Lemma 6.4 implies that $D$ must have parameters
@@
Observe that if $p=4t^{2}+9$, then $p-1$ can be of the form $p-1=4m$, where $m$ is odd and square-free. For example, $p=3373=4\cdot 29^{2}+9=4\cdot 3\cdot 281+1$. It would be interesting to study schurity of $D_{2p}$ for all primes $p$ such that $p\equiv 1\bmod 4$.
8.3. Schurity of Frobenius groups
Recall that a Frobenius group $C_{p}\rtimes C_{p-1}$ is not Schur for every prime $p\geq 7$ by Corollary 5.7. It seems interesting to ask which Frobenius groups are Schur. Denote the set of all integers of one of the forms
$p^{k},~{}pq^{k},2pq^{k},~{}pqr,~{}2pqr,$
where $p$, $q$, and $r$ are primes and $k\geq 1$, by $\mathcal{P}$ and the set of all odd integers from $\mathcal{P}$ by $\mathcal{P}_{0}$. The proposition below provides several restrictions on the structure of a Frobenius Schur group.
Proposition 8.3.
Let $G$ be a Frobenius Schur group. Then $G$ is isomorphic to one of the following groups:
  1. (1)
    $C_{n}\rtimes C_{m}$, where $n\in\mathcal{P}_{0}$ and $m\in\mathcal{P}$;
  2. (2)
    $(E_{4}\times C_{p^{k}})\rtimes C_{3}$, $(E_{4}\times C_{pq})\rtimes C_{3}$, where $p$ and $q$ are odd primes and $k\geq 1$;
  3. (3)
    $E_{4}\rtimes C_{3}$, $E_{27}\rtimes C_{13}$, $E_{32}\rtimes C_{31}$.
Proof.
We start the proof with the lemma below which follows from computational results [44] and [32, Lemma 3.7].
Lemma 8.4.
The Frobenius groups $E_{8}\rtimes C_{7}$, $E_{9}\rtimes C_{2}$, $E_{16}\rtimes C_{3}$, $E_{16}\rtimes C_{5}$, $E_{27}\rtimes C_{2}$ are not Schur.
Let $G=N\rtimes A$ be a Frobenius Schur group, where $N$ and $A$ are the Frobenius kernel and complement of $G$, respectively. From [32, Theorem 1.3] it follows that $G$ is metabelian and hence the derived subgroup $G^{\prime}$ is abelian. Since $G$ is Frobenius, $G^{\prime}=N\rtimes A^{\prime}$. If $A^{\prime}$ is nontrivial, then $G^{\prime}$ is also Frobenius and, in particular, nonabelian. So $A^{\prime}$ is trivial and consequently $N$ and $A$ are abelian. By Lemma 2.2, the groups $N$ and $A$ are Schur. Due to [20, Corollary 6.17], $A$ is cyclic. Therefore [14, Theorem 1.1] yields the following
Lemma 8.5.
The Frobenius complement $A$ is a cyclic group of order $n\in\mathcal{P}$.
Observe that $N$ does not have a characteristic subgroup of order $2$. Together with [32, Theorem 3.3], this implies the following
Lemma 8.6.
The Frobenius kernel $N$ is isomorphic to one of the groups below:
  1. (1)
    a cyclic group of order $n\in\mathcal{P}_{0}$;
  2. (2)
    $E_{n}$, where $n\in\{4,8,9,16,27,32\}$;
  3. (3)
    $C_{3}\times C_{3^{k}}$, $E_{9}\times C_{p}$, where $k\geq 2$ and $p\neq 2$ is a prime;
  4. (4)
    $E_{4}\times C_{p^{k}}$, $E_{4}\times C_{pq}$, $E_{16}\times C_{p}$, where $k\geq 1$ and $p\neq 2$ and $q$ are primes.
If Statement $(1)$ of Lemma 8.6 holds for $N$, then Statement $(1)$ of Proposition 8.3 holds for $G$. Further, we assume that $N$ is noncyclic and hence one of Statements $(2)$-$(4)$ of Lemma 8.6 holds for $N$.
Suppose that $|N|$ is even. Then $P=\operatorname{Syl}_{2}(N)$ is nontrivial and $|A|$ is odd. From Lemma 8.6 it follows that $P$ is isomorphic to one of the groups
$E_{4},~{}E_{8},~{}E_{16},~{}E_{32}.$
Since $P$ is characteristic in $N$, the group $H=P\rtimes A$ is Frobenius. By Lemma 2.2, the group $H$ and each of its subgroups are Schur. Clearly, $|A|$ divides $|P|-1$. If $P\cong E_{8}$, then $H\cong E_{8}\rtimes C_{7}$, a contradiction to Lemma 8.4. If $P\cong E_{16}$, then $H$ is isomorphic to one of the groups $E_{16}\rtimes C_{3}$, $E_{16}\rtimes C_{5}$, $E_{16}\rtimes C_{15}$. In all cases, $H$ has a subgroup isomorphic to $E_{16}\rtimes C_{3}$ or $E_{16}\rtimes C_{5}$, a contradiction to Lemma 8.4.
If $P\cong E_{4}$, then $N$ is isomorphic to one of the groups $E_{4}$, $E_{4}\times C_{p^{k}}$, $E_{4}\times C_{pq}$ by Lemma 8.6 and $A\cong C_{3}$. So Statement $(2)$ or $(3)$ of Proposition 8.3 holds for $G$. If $P\cong E_{32}$, then $N\cong E_{32}$ by Lemma 8.6 and $A\cong C_{31}$. In this case, Statement $(3)$ of Proposition 8.3 holds for $G$.
Now suppose that $|N|$ is odd. Then Lemma 8.6 yields $N$ has a characteristic subgroup isomorphic to $E_{9}$ or $N\cong E_{27}$. In the first case, $G$ has a subgroup isomorphic to $E_{9}\rtimes C_{2}$ which is not Schur by Lemma 8.4, a contradiction to Lemma 2.2. In the second one, $A$ is isomorphic to one of the groups $C_{2}$, $C_{13}$, $C_{26}$ because $|A|$ divides $|N|-1=26$. If $A\cong C_{2}$ or $A\cong C_{26}$, then $G$ has a subgroup isomorphic to $E_{27}\rtimes C_{2}$ which is not Schur by Lemma 8.4, a contradiction to Lemma 2.2. If $A\cong C_{13}$, then $G\cong E_{27}\rtimes C_{13}$ and Statement $(3)$ of Proposition 8.3 holds for $G$. ∎
It seems reasonable to ask about schurity of groups from Proposition 8.3. Computer calculations [44] imply that all Frobenius groups of order at most $62$ satisfying the conditions from Proposition 8.3 are Schur except for the Frobenius group $C_{19}\rtimes C_{3}$ which is not Schur. A unique nonschurian $S$-ring over this group is a primitive $S$-ring of rank $3$ arising from a strongly regular Cayley graph which is a block graph of a Steiner triple system. A construction of this $S$-ring is described in detail in [31].